# Improving Image Captioning Descriptiveness by Ranking and LLM-based Fusion

### v1

@@ Abstract
State-of-The-Art (SoTA) image captioning models often rely on the Microsoft COCO (MS-COCO) dataset for training. This dataset contains annotations provided by human annotators, who typically produce captions averaging around ten tokens. However, this constraint presents a challenge in effectively capturing complex scenes and conveying detailed information. Furthermore, captioning models tend to exhibit bias towards the “average” caption, which captures only the more general aspects. What would happen if we were able to automatically generate longer captions, thereby making them more detailed? Would these captions, evaluated by humans, be more or less representative of the image content compared to the original MS-COCO captions? In this paper, we present a novel approach to address previous challenges by showcasing how captions generated from different SoTA models can be effectively fused, resulting in richer captions. Our proposed method leverages existing models from the literature, eliminating the need for additional training. Instead, it utilizes an image-text based metric to rank the captions generated by SoTA models for a given image. Subsequently, the top two captions are fused using a Large Language Model (LLM). Experimental results demonstrate the effectiveness of our approach, as the captions generated by our model exhibit higher consistency with human judgment when evaluated on the MS-COCO test set. By combining the strengths of various SoTA models, our method enhances the quality and appeal of image captions, bridging the gap between automated systems and the rich, informative nature of human-generated descriptions. This advance opens up new possibilities for generating captions that are more suitable for the training of both vision-language and captioning models [^1] Code and data are available at https://drive.google.com/drive/folders/1tXbJyvwrEF-a-t-NZbt2xjoKaiGGeaz1?usp=sharing..
@@ 1 Introduction
Image captioning endeavors to produce organic and human-like descriptions corresponding to a given image. This task holds substantial significance in numerous practical applications, including human-computer interaction and multi-modal recommendation systems. Consequently, it has garnered considerable research interest in recent times. Notably, several State-of-The-Art (SoTA) models have emerged Alayrac et al. (2022); Hu et al. (2022a); Wang et al. (2022b); Yu et al. (2022), demonstrating highly encouraging outcomes when evaluated against widely used metrics such as BLEU Papineni et al. (2002), CIDEr Vedantam et al. (2015), METEOR Banerjee and Lavie (2005), and SPICE Anderson et al. (2016). These models are typically trained on the Microsoft COCO (MS-COCO) dataset Lin et al. (2014), which contains captions provided by human annotators. MS-COCO has certain limitations: (i) consists of images depicting common objects and scenes, but lacks images of rare or complex events, such as disasters, sports, or art; (ii) captions and queries associated with images often tend to be simplistic, repetitive or inaccurate, which does not reflect the use of natural language and user expectations. Therefore, there is a need for more realistic, large-scale and diverse datasets that can capture the variety and richness of visual and linguistic information in the real-world. The collection of high-quality annotated image-text pair datasets is both time and cost prohibitive. Recently, some vision-language alignment approaches have exploited SoTA captioners for automatically describing the content of images collected from the web Li et al. (2022, 2021a); Wang et al. (2022c). However, the alt-texts often do not accurately describe the visual content of the images, making them a noisy signal that is suboptimal for learning vision-language alignment.
In this paper we propose a captioning model for the creation of high-quality and high-descriptive image-text pair datasets without requiring human intervention. Given an image, our novel training-free model generates captions leveraging several SoTA captioners, then uses a text-image based metric to rank them. The first two captions are finally merged using a Large Language Model (LLM) to generate the final caption. The proposed model will also allow to answer the following Research Questions (RQs):
@@
    Would the generated captions be more representative of the image content according to human judgement?
@@
    Can current image caption models generate richer and more descriptive captions if they are trained using ground-truth data that exhibits these characteristics?
Experimental results demonstrate the effectiveness of our model, as the captions generated by our model exhibit higher consistency with human judgment when evaluated on the MS-COCO test set. By combining the strengths of various SoTA models, our method enhances the quality and appeal of image captions, bridging the gap between automated systems and the rich, informative nature of human-generated descriptions.
@@ 2 Related works
Since the advent of deep learning as the predominant approach, there has been significant advancement in image captioning Anderson et al. (2018); Donahue et al. (2015); Hu et al. (2022a); Huang et al. (2019). Transformer-based architectures have gained particular prominence in recent times Li et al. (2020); Shen et al. (2022). Many research papers follow the strategy of initially pretraining large vision-language models and then adapting them for specific tasks such as captioning Li et al. (2023, 2021b); Wang et al. (2022c). Present efforts mainly revolve around further scaling these pretraining-based methods Alayrac et al. (2022); Hu et al. (2022b); Wang et al. (2022a); Yu et al. (2022) and integrating various vision-language tasks during the pretraining stage Cho et al. (2021); Wang et al. (2022b).
@@ 3 Proposed model
Figure 1 shows an overview of the proposed image captioning model. Given $N$ SoTA image captioning models and an image $I$, it is possible to generate $N$ different captions $\{T_{1},...,T_{N}\}$ that describe the content of the image. Generated captions are then ranked according to their alignment to the image content through an Image-Text Matching (ITM) method. Finally, the meaning of the top-2 ranked captions is fused by a LLM to generate a new caption.
@@ 3.1 Caption generation
We have carefully selected five cutting-edge image caption model that excel in terms of performance, reproducibility, scalability, and popularity. Our decision to choose five models aligns with the number of ground-truth captions provided by MS-COCO. However, it is important to note that this number can be easily scaled up to accommodate a higher number of models, albeit at an increased computational cost. The selected models include Bootstrapping Language-Image Pre-training 2 (BLIP-2) Li et al. (2023), ExpansionNet v2 (referred to as ExpNet-v2) Hu et al. (2022a), Generative Image-to-text Transformer (GIT) Wang et al. (2022a), One For All (OFA) Wang et al. (2022b), and ViT-GPT2 NLP Connect (2022). For models that have multiple variants, the largest publicly available variant is chosen to ensure greater effectiveness. In particular for BLIP-2 the ViT-g OPT2.7B is used, GITLarge with the ViT-L/14 from CLIP Radford et al. (2021) image encoder is selected, finally OFAHuge exploiting a ResNet-152 as backbone and a Transformer architecture. We point out that the proposed approach is not limited to the number or models that have been chosen for current experiments. Table 1 presents the key characteristics of the selected models. Additional details can be found in the Appendix A.1.
@@
BLIP-2 (2023) Li et al. (2023) ViT-g OPT2.7B 3.8B
ExpansionNet v2 (2022) Hu et al. (2022a) SwinT-L FeedForwardNet 236M
GIT (2022) Wang et al. (2022a) ViT-L/14 Transformer 394M
OFA (2022) Wang et al. (2022b) ResNet152+Transformer BPE 930M
ViT-GPT2 (2022) NLP Connect (2022) ViT-B/16 GPT2 302M
@@
Image
BLIP-2
ExpNet-v2
GIT
OFA
ViT-GPT2
@@ 3.2 Caption ranking
The goal of this step is to select two captions from the entire set of generated captions $\{T_{1},...,T_{N}\}$. We first define the overall matching score $o_{i}$ of each caption. The captions generated by the chosen models can be evaluated in terms of fidelity, i.e., how much the generated caption is related to the input image without distortion; of adequacy, that is, how much image gist it conveys; fluency, how fluent and beautiful are the language and words Wang et al. (2021). Since the above three aspects are related to both image and text modalities, we use a multi-modal method to estimate the match between each generated caption and the input image. Specifically, we exploit BLIP for Image-Text Retrieval (ITR) Li et al. (2022). The pretrained BLIP is finetuned for ITR by minimizing the Image-Text Matching (ITM) and Image-Text Contrastive (ITC) loss functions. An input image $I$ is encoded by the image encoder into a sequence of embeddings: $\{\mathbf{v}_{\mathrm{cls}},\mathbf{v}_{1},...,\mathbf{v}_{L}\}$, where $\mathbf{v}_{\mathrm{cls}}$ is the embedding of the [CLS] token. The text encoder transforms each generated caption $T_{i}$ into a sequence of embeddings $\{\mathbf{w}_{\mathrm{cls}},\mathbf{w}_{1},...,\mathbf{w}_{M}\}$.
@@
Matching probability. The image features are fused with the text features through cross attention at each layer of the Image-grounded text encoder. A Fully-Connected (FC) layer followed by softmax predict a two-class probability $\mathbf{p}\in\mathbb{R}^{2}$ indicating the image-text matching probability.
@@
Model
Caption
Matching probability Cosine similarity BLIPScore
@@ 3.3 Caption fusion
The first two captions of the rank, namely $T_{\pi_{1}}$ and $T_{\pi_{2}}$, are selected to generate a new caption. The GPT-3 model, also known as Davinci, developed by OpenAI, is utilized for captioning fusion Brown et al. (2020). Davinci stands out as one of the largest and most proficient LLMs currently available. It undergoes training on an extensive corpus of textual data using the unsupervised Transformer learning mechanism, a pre-built learning technique. As a result, the model autonomously learns natural language structures, syntax, rules, and semantics, achieving text generation capabilities comparable to human-like proficiency. In addition to text generation, Davinci demonstrates the ability to analyze sentence syntax and semantics to extract relevant information and generate answers to queries or questions to the best of its abilities.
Hence, to merge the two best generated captions, the following question is prompted to Davinci: “Combine the meaning of these 2 sentences into 1 sentence, considering the semantic meaning and the syntactic meaning. The sentences are: caption1; caption2. These sentences describe an image, I want to get the best caption of the image, using the information in these two sentences.” Here, caption1 and caption2 are $T_{\pi_{1}}$ and $T_{\pi_{2}}$, namely the two captions with the highest BLIPScore. The question to ask Davinci was engineered and the different considerations that lead to the final version can be found in the Appendix A.2.
Figure 4 presents an image from the MS-COCO test set accompanied by the five ground-truth captions, the five captions generated by the SoTA models, and the caption generated by our model. The notable observation is that our generated caption is not only more detailed but also more articulate than competitors. This can be attributed to the advanced generative capabilities of Davinci, which utilizes a larger and more extensive vocabulary compared to conventional image captioning models. The utilization of the Davinci model as a text generator significantly enhances the grammar, syntax, and lexical diversity of the fused caption, thereby producing a remarkably human-like description. For additional qualitative results, please refer to Appendix B.2.

### v3

@@ Abstract
State-of-The-Art (SoTA) image captioning models are often trained on the MicroSoft Common Objects in Context (MS-COCO) dataset, which contains human-annotated captions with an average length of approximately ten tokens. Although effective for general scene understanding, these short captions often fail to capture complex scenes and convey detailed information. Moreover, captioning models tend to exhibit bias towards the “average” caption, which captures only the more general aspects, thus overlooking finer details. In this paper, we present a novel approach to generate richer and more informative image captions by combining the captions generated from different SoTA captioning models. Our proposed method requires no additional model training: given an image, it leverages pre-trained models from the literature to generate the initial captions, and then ranks them using a newly introduced image-text-based metric, which we name BLIPScore. Subsequently, the top two captions are fused using a Large Language Model (LLM) to produce the final, more detailed description. Experimental results on the MS-COCO and Flickr30k test sets demonstrate the effectiveness of our approach in terms of caption-image alignment and hallucination reduction according to the ALOHa, CAPTURE, and Polos metrics. A subjective study lends additional support to these results, suggesting that the captions produced by our model are generally perceived as more consistent with human judgment. By combining the strengths of diverse SoTA models, our method enhances the quality and appeal of image captions, bridging the gap between automated systems and the rich and informative nature of human-generated descriptions. This advance enables the generation of more suitable captions for the training of both vision-language and captioning models.
keywords:
Image captioning, Large Language Model, Image-text matching, Descriptiveness.
@@ 1 Introduction
Figure 1: Comparison between ground-truth captions (Original) and enriched captions generated by our method (Ours) for sample images from the COCO dataset.
Image captioning endeavors to produce organic and human-like descriptions corresponding to a given image. This task holds substantial significance in numerous practical applications, including human-computer interaction and multi-modal recommendation systems. Consequently, it has garnered considerable research interest in recent times. Notably, several State-of-The-Art (SoTA) models have emerged [1, 24, 19], demonstrating highly encouraging outcomes when evaluated against widely used metrics such as BLEU [32], CIDEr [42], METEOR [5], SPICE [2]. These models are typically trained on the Microsoft COCO (MS-COCO) dataset [27] and the Flickr30k dataset [49], which contain captions provided by human annotators. Both MS-COCO and Flickr30k have certain limitations: (i) they mainly consist of images depicting common objects and scenes, but lack images of rare or complex events, such as disasters, sports, or art; (ii) the captions and queries associated with the images often tend to be simplistic, repetitive, or inaccurate, which does not reflect the use of natural language and user expectations, as illustrated in Figure 1. Therefore, there is a need for more realistic, large-scale, and diverse datasets that can capture the variety and richness of visual and linguistic information in the real world. The collection of high-quality annotated image-text pair datasets is both time and cost-prohibitive. Recently, some vision-language alignment approaches have exploited SoTA captioning models to automatically describe the content of images collected from the web [22, 21, 47]. However, the alt-texts often do not accurately describe the visual content of the images, making them a noisy signal that is suboptimal for learning vision-language alignment.
In this paper, we propose a captioning model designed to create high-quality and highly descriptive image-text pair datasets without requiring human intervention. Given an image, the model generates candidate captions leveraging several SoTA captioners and ranks them using a text–image–based metric. The top-ranked captions are then merged using a Large Language Model (LLM) to produce the final output. We evaluate our approach on the MS-COCO and Flickr30k datasets and adopt a broad set of evaluation metrics, including BLEU, CIDEr, METEOR, SPICE, and more recent metrics such as ALOHa [33], CAPTURE [11], and Polos [43], to comprehensively assess factuality, image–text alignment, and semantic consistency. In addition to these automatic metrics, we also perform a subjective human evaluation to measure user preference and perceived caption quality. Furthermore, we experiment with different LLM backbones for the fusion stage (such as Davinci, Mixtral-8x7B-Instruct, and Llama-3.3-70B-Instruct) and conduct ablation studies to analyze the impact of each component of the pipeline. Finally, we demonstrate that training an existing captioning model on our fused captions as ground-truth leads to consistent improvements across multiple metrics, indicating that the captions generated by our method can also benefit downstream model learning and generalization. The proposed model will enable answering the following Research Questions (RQs):
@@
    Would the generated captions be more representative of the image content according to human judgment?
@@
    Can training existing captioning models on the fused captions produced by our method improve their performance and generalization ability?
Experimental results demonstrate the effectiveness of our model, as the captions generated by our model exhibit greater consistency with human judgment when evaluated on the MS-COCO and Flickr30k test sets. By combining the strengths of various SoTA models, our method enhances the quality and appeal of image captions, bridging the gap between automated systems and the rich and informative nature of human-generated descriptions.
@@ 2 Related works
In this section, we review prior work in image captioning, focusing on SoTA models, the use of large-scale web data, and recent approaches to caption enrichment.
2.1 Image captioning
Since the advent of deep learning as the predominant approach, there has been a significant advancement in image captioning [3, 10, 15, 17]. Transformer-based architectures have gained particular prominence in recent times [26, 38]. ExpansionNet v2 [15] utilizes a Swin Transformer encoder and an expansion-based decoder for improved caption generation, incorporating dynamic expansion and multi-head attention mechanisms. ViT-GPT2 [30] combines a ViT encoder and a GPT-2 decoder for a straightforward and effective multimodal approach, leveraging the strengths of pretrained unimodal models. Many research papers follow the strategy of initially pretraining large vision-language models and then adapting them for specific tasks such as captioning [23, 25, 47]. BLIP-2 [23] integrates pretrained Vision-Language Processing (VLP) and NLP models using a Querying Transformer (Q-Former) to bridge the image encoder and LLM. It employs a two-phase pretraining strategy for representation alignment and generative learning. Present efforts mainly revolve around further scaling these pretraining-based methods [1, 16, 44, 50] and integrating various vision-language tasks during the pretraining stage [9, 45]. OFA [45] is a task-agnostic model unifying image and text representations using Transformers, fine-tuned with cross-entropy and CIDEr optimization for image captioning. GIT [44] adopts a vision transformer-based image encoder and a Transformer text decoder, pretrained on extensive image-text datasets for mapping visual inputs to text descriptions. I-Tuning [28] is a lightweight image captioning model that introduces a novel cross-attention module to bridge a frozen GPT-2 language decoder with a CLIP-ViT vision encoder. CA-Captioner [48] enhances sentence hierarchy using a concentrated attention mechanism consisting of three components: Head Absolute Positional Encoding captures spatial relationships, Learnable Sparse Mechanism filters noise and emphasizes key objects, and Local Feature Enhancement integrates local detail features. EVCap [24] is a retrieval-augmented image captioning model that enriches a frozen LLM with object names from an external visual–name memory. It uses a lightweight attentive fusion module to combine retrieved names and visual features, enabling open-world comprehension with just 3.97M trainable parameters and no fine-tuning on out-of-domain data. MSRM [13] generates captions with rich scene details and accurate relationships through three key innovations: a Semantic Cue Miner to extract dynamic semantic cues, a Semantic Mapper to establish fine-grained relational interactions, and an Adaptive Bridging Decoder to dynamically fuse multi-granularity features. VIPCap [19] is a retrieval-based visual prompt for lightweight image captioning. It transforms text retrieved for an image into semantic features using a learnable Gaussian distribution and aligns them with visual features to generate visual prompts.
Previous methods generate captions from scratch by optimizing supervised loss on curated datasets. In contrast, our method does not require any training or fine-tuning. It leverages multiple pretrained captioners as black-box generators to create candidate captions, then fuses the most relevant ones using an LLM. This strategy allows us to integrate diverse semantic perspectives without modifying or retraining the underlying captioning architectures.
2.2 Image captioning datasets
Image-caption datasets generally fall into two main categories: human-annotated datasets like MS-COCO [27] and Flickr30k [49], and web-crawled datasets such as CC, CC12M, and SBU Captions [7, 37, 31]. These datasets collectively form the foundation for VLP and are subsequently fine-tuned for downstream tasks. Typically, human-annotated datasets are smaller but have significantly less noise compared to web-crawled datasets. However, both categories are characterized by relatively short and concise captions.
Data-centric AI emphasizes improving data quality to enhance model performance, focusing on curating, labeling, and cleaning datasets [41], as well as automating these processes [52]. This perspective suggests that high-quality data can enable even simple algorithms to achieve remarkable results. Addressing the need for enriched image-text datasets, Shi et al[39] proposed using natural language inference to merge multiple ground-truth captions into a single, more comprehensive one. However, this method is limited to datasets with multiple ground-truth captions and is not applicable to large-scale datasets with only a single caption per image (e.g., CC, CC12M, and SBU).
2.3 Image captioning enhancement
Recent approaches aim to generate enriched image captions by merging multiple sources or incorporating additional semantic information derived from the visual content. LaCLIP [12] uses a LLM to rewrite raw captions, but its performance is hindered by issues such as hallucinations—often due to limited visual grounding and low-quality input captions. FuseCap [36] enhances captions using visual experts like object detectors. VeCLIP [20] uses an LLM to combine raw and synthetic captions, but it depends directly on a pre-existing LLM for inference and lacks explicit guidance, such as leveraging world knowledge from raw captions or syntactic cues from synthetic ones. CapsFusion [51] addresses this by fine-tuning an open-source LLM on data generated by ChatGPT and incorporating detailed instructions, which help the model make more informed decisions during fusion. The Visual Fact Checker [14] focuses on reducing hallucinations in extended captions by integrating outputs from two multimodal captioners. It uses object detection for fact-checking and an LLM to verify and consolidate the results into a coherent final caption. QAC [29] enhances caption detail by combining a Questioner (which queries objects, spatial relationships, and world knowledge), ChatGPT (for comprehensive responses), and an Answerer (which grounds responses in visual data), with the final caption synthesized by a dedicated Captioner. Compared to previous methods, our approach: (i) requires no additional training or supervision; (ii) employs pretrained captioners—trained on large-scale web data—to generate diverse, high-quality descriptions; (iii) uses an automatic ranking mechanism (see Section 3.2) to select the most semantically aligned candidate captions; (iv) performs prompt-based fusion via an off-the-shelf LLM, without relying on retrieval modules, object detectors, or fine-tuning.
@@ 3 Proposed model
Figure 2 shows an overview of the proposed image captioning model. Given $N$ SoTA image captioning models and an image $I$, it is possible to generate $N$ different captions $\{T_{1},...,T_{N}\}$ that describe the content of the image. The generated captions are then ranked according to their alignment with the image content using an Image-Text Matching (ITM) method. Finally, the meaning of the top-2 ranked captions is fused by an LLM to generate a new caption.
@@ 3.1 Caption generation
We have carefully selected five cutting-edge image caption models that excel in terms of performance, reproducibility, scalability, and popularity. We selected five models to align with the number of ground-truth captions provided by MS-COCO. However, it is important to note that this number can be easily scaled up to accommodate a larger number of models, albeit at an increased computational cost. The selected models include Bootstrapping Language-Image Pre-training 2 (BLIP-2) [23], ExpansionNet v2 (referred to as ExpNet-v2) [15], Generative Image-to-text Transformer (GIT) [44], One For All (OFA) [45], and ViT-GPT2 [30]. For models that have multiple variants, the largest publicly available variant is chosen to ensure greater effectiveness. In particular for BLIP-2 the ViT-g OPT2.7B is used, GITLarge with the ViT-L/14 from CLIP [34] image encoder is selected, finally OFAHuge which exploits a ResNet-152 as backbone and a Transformer architecture. We point out that the proposed approach is not limited to the number or models that have been chosen for current experiments. Table 1 summarizes the key characteristics of the selected models.
@@
BLIP-2 [23] ViT-g OPT2.7B 3.8B
ExpansionNet v2 [15] SwinT-L FeedForwardNet 236M
GIT [44] ViT-L/14 Transformer 394M
OFA [45] ResNet152+Transformer BPE 930M
ViT-GPT2 [30] ViT-B/16 GPT2 302M
@@
Image BLIP-2 ExpNet-v2 GIT OFA ViT-GPT2
@@ 3.2 Caption ranking
The goal of this step is to select two captions from the entire set of generated captions $\{T_{1},...,T_{N}\}$. We first define the overall matching score $o_{i}$ of each caption. The captions generated by the chosen models can be evaluated in terms of fidelity, i.e., how much the generated caption is related to the input image without distortion; of adequacy, that is, how much image gist it conveys; fluency, how fluent and natural are the language and word choice [46]. While all three aspects are taken into account in the evaluation process, only fidelity and adequacy are implicitly modeled by the designed metric, as they directly pertain to the semantic alignment between image and text. To estimate this alignment, we adopt a multi-modal approach using BLIP for Image-Text Retrieval (ITR) [22]. The pretrained BLIP is finetuned for ITR by minimizing the Image-Text Matching (ITM) and Image-Text Contrastive (ITC) loss functions. An input image $I$ is encoded by the image encoder into a sequence of embeddings: $\{\mathbf{v}_{\mathrm{cls}},\mathbf{v}_{1},...,\mathbf{v}_{L}\}$, where $\mathbf{v}_{\mathrm{cls}}$ is the embedding of the [CLS] token. The text encoder transforms each generated caption $T_{i}$ into a sequence of embeddings $\{\mathbf{w}_{\mathrm{cls}},\mathbf{w}_{1},...,\mathbf{w}_{M}\}$.
@@
Matching probability. The image features are fused with the text features through cross attention at each layer of the Image-grounded text encoder. A Fully-Connected (FC) layer followed by softmax predicts a two-class probability $\mathbf{p}\in\mathbb{R}^{2}$ indicating the image-text matching probability.
@@
Model Caption
Matching
probability
Cosine
similarity
BLIPScore
@@ 3.3 Caption fusion
The first two captions of the rank, namely $T_{\pi_{1}}$ and $T_{\pi_{2}}$, are selected to generate a new caption. The GPT-3 model, also known as Davinci, developed by OpenAI, is utilized for captioning fusion [6]. Davinci stands out as one of the largest and most proficient LLMs currently available. It undergoes training on an extensive corpus of textual data using the unsupervised Transformer learning mechanism, a pre-built learning technique. As a result, the model autonomously learns natural language structures, syntax, rules, and semantics, achieving text generation capabilities comparable to human-like proficiency. In addition to text generation, Davinci demonstrates the ability to analyze sentence syntax and semantics to extract relevant information and generate answers to queries or questions to the best of its abilities.
Hence, to merge the two best generated captions, the following question is prompted to Davinci: “Combine the meaning of these 2 sentences into 1 sentence, considering the semantic meaning and the syntactic meaning. The sentences are: caption1; caption2. These sentences describe an image, I want to get the best caption of the image, using the information in these two sentences.” Here, caption1 and caption2 are $T_{\pi_{1}}$ and $T_{\pi_{2}}$, namely the two captions with the highest BLIPScore. The question to ask Davinci was engineered, and the different considerations that led to the final version can be found in the Ablation study section.
Figure 5 presents an image from the MS-COCO test set accompanied by the five ground-truth captions, the five captions generated by the SoTA models, and the caption generated by our model. The notable observation is that our generated caption is not only more detailed, but also more articulated than competitors. This can be attributed to the advanced generative capabilities of Davinci, which utilizes a larger and more extensive vocabulary compared to conventional image captioning models. The utilization of the Davinci model as a text generator significantly enhances the grammar, syntax, and lexical diversity of the fused caption, thereby producing a remarkably human-like description.